Calculate the molar solubility of Mg(OH)2 in a solution that is basic with a pH of 12.62.

Ksp = [Mg2+][OH]2 = 5.6 × 10–12
Answer #1

Mg (OH)2 (s)Mg2*(aq)+20H (ag) pH-12.62 pOH -14- pH =14-1262 OH-10-P0H 1.38 = 10- 0.0417M Molar solubility of Mg(OH)2 in 0.041